OpenStudy (anonymous):

PEMDAS, NEED HELP! 18 divided by 9 * 2 + 3, please show me how you do it! Thank you!

OpenStudy (anonymous):

18/9=2

OpenStudy (anonymous):

2*2=4

OpenStudy (anonymous):

4+3= 7

OpenStudy (anonymous):

that is following Pemdas